### v1

# Smart Expert System: Large Language Models as Text Classifiers

Abstract
Text classification is a fundamental task in Natural Language Processing (NLP), and the advent of Large Language Models (LLMs) has revolutionized the field. This paper introduces the Smart Expert System, a novel approach that leverages LLMs as text classifiers. The system simplifies the traditional text classification workflow, eliminating the need for extensive preprocessing and domain expertise. The performance of several LLMs, machine learning (ML) algorithms, and neural network (NN) based structures is evaluated on four datasets. Results demonstrate that certain LLMs surpass traditional methods in sentiment analysis, spam SMS detection and multi-label classification. Furthermore, it is shown that the system’s performance can be further enhanced through few-shot or fine-tuning strategies, making the fine-tuned model the top performer across all datasets. Source code and datasets are available in this GitHub repository: https://github.com/yeyimilk/llm-zero-shot-classifiers.
Index Terms:
Large Language Models, Text Classification, Natural Language Processing, Smart Expert Systems, Few-Shot Learning, Fine-Tuning, Chat GPT-4, Llama3.
Figure 1: Traditional text classification flow
Figure 2: LLMs’ zero-shot text classification simple flow
I Introduction
Text classification is a key task in NLP, widely used in areas like sentiment analysis Liu (2022), topic labeling Chen et al. (2020), question answering Minaee et al. (2021), and dialog act classification Qin et al. (2020). Traditional ML methods, such as logistic regression, Naive Bayes, and support vector machines, are common but need a lot of labeled data and can only handle known classes Sarker (2021); Wang et al. (2019). Deep learning (DL) methods, including deep neural networks, recurrent neural networks, and convolutional neural networks, have surpassed traditional ML algorithms by capturing complex data relationships LeCun et al. (2015); Sutskever et al. (2011). However, DL methods also require large labeled datasets and face challenges in adapting to new tasks without significant retraining.
LLMs, especially those based on the Transformer architecture like PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023), GPT Radford et al. (2018), and BERT Devlin et al. (2018), have revolutionized NLP. With hundreds of billions to trillions of parameters, these models are pre-trained on massive amounts of text data. Compared to traditional DL methods, LLMs are much larger and exhibit stronger language understanding and generation capabilities. They learn rich language representations through extensive pretraining, enabling them to adapt quickly to various downstream tasks. For instance, Přibáň et al. (2024) used LLMs for cross-lingual sentiment analysis. Caruccio et al. (2024) explored using LLMs for intelligent diagnoses. DePalma et al. (2024) examined LLMs’ code refactoring capabilities. Additionally, Boubker (2024) and Oh et al. (2024) evaluated and enhanced student learning performance using LLMs.
Although LLMs have been widely applied and researched for various tasks, their potential for handling text classification tasks remains to be further explored. Recent studies indicate that while LLMs perform excellently in syntactic analysis, there are still numerous challenges in achieving efficient and accurate text classification in practical applications. The studies by Bețianu et al. (2024) and Møller et al. (2024) demonstrate the application of LLMs in domain adaptation and synthetic data augmentation, revealing both the strengths and limitations of these models in handling rare categories or domain shifts. However, the development of LLMs offers new research opportunities to simplify and optimize the text classification process.
Traditional ML and NN approaches to text classification involve a multi-stage pipeline that can be both complex and resource-intensive. These stages typically include feature extraction, dimensionality reduction, classifier selection, and model evaluation Kowsari et al. (2019). Figure 1 illustrates this conventional workflow, which starts with data collection followed by laborious preprocessing steps such as tokenization, punctuation removal, stop word filtering, stemming, and lemmatization. Subsequently, feature extraction techniques like bag-of-words, word embeddings, or TF-IDF are applied before training a classifier to produce the final classification model.
However, this traditional approach presents several challenges. Firstly, it requires significant domain expertise to effectively preprocess the data and engineer features that capture the relevant information for the classification task. Secondly, each step in the process must be carefully tuned to optimize performance—a time-consuming endeavor that often involves trial-and-error experimentation. Lastly, these methods may not generalize well across different datasets or languages without substantial reconfiguration.
In contrast, LLMs offer a much easier way to do text classification, as depicted in Figure 2, that it just includes three main blocks: data collection, feeding data directly into LLMs, and obtaining classification results from LLMs. This simplified flow eliminates the need for explicit feature extraction or dimensionality reduction since LLMs inherently encode rich linguistic features through their deep contextual representations. Moreover, because they have been pre-trained on diverse datasets, they require minimal—if any—additional training specific to a given task or domain.
In light of these considerations, we propose a novel approach that capitalizes on the strengths of LLMs for text classification tasks. Our approach is encapsulated in the Smart Expert System, which utilizes LLMs as Text Classifiers, as depicted in Figure 3. This system is designed to leverage the advanced capabilities of LLMs, offering a more streamlined and efficient method for classifying text without the extensive preprocessing and feature engineering required by traditional ML/NN methods. Introducing our Smart Expert System represents a significant shift towards more intelligent and adaptable NLP tools. By integrating LLMs into the core of text classification workflows, we aim to reduce complexity and enhance performance across various domains and applications.
Our main contributions are as follows:
  • This paper proposes a text classification system using LLMs to replace traditional text classifiers. This system simplifies the conventional text classification process, lowering technical barriers and eliminating the need for domain experts to perform complex preprocessing and algorithm design. This approach is crucial for rapid deployment and scalable applications, especially for small businesses that do not have deep expertise in ML or DL.
  • We introduce a new performance evaluation metric, the Uncertainty/Error Rate (U/E rate). This metric supplements traditional accuracy and F1 scores, providing a more comprehensive evaluation of a model’s performance under unknown or uncertain conditions and emphasizing the LLMs’ reliability in real-world applications.
  • We compare the performance of LLMs with traditional ML and NN models across multiple datasets. The results show that LLMs, after learning from a small number of samples or fine-tuning, outperform in various text classification tasks. This finding confirms the versatility and efficiency of LLMs.
Figure 3: Framework of our Expert System. The steps of the framework can be included as (1) collect data from the data source to establish the domain database; (2) send domain-specific data to the pre-trained LLM model, like GPT-4, Llama-3 and so on; (3) using a few domain-specific data to do fine-tuning or few-shot learning; (4) apply the fine-tuning or few-shot learning to the pre-trained LLM model; (5) (optional) utilize domain knowledge expert to set up the prompts to elevate LLM performance; (6) apply prompts in the pre-trained model; (7) evaluate the whole expert system’s performance; (8) non-expert users query tasks through user interface to the expert system; (Tasks may include classification, sentiment analysis, prediction, recommendation and so on. In this paper, we take the classification and sentiment analysis as examples.) (9) LLM API interacts with User interface and the pre-trained LLM model, advising on the user interface.
II Background and Related Work
II-A Traditional Text Classification Approaches
Text classification has evolved through several methodologies, each with unique advantages and limitations. Early rule-based methods like decision trees and expert systems, such as C4.5 Quinlan (2014) and MYCIN Shortliffe (2012), were straightforward but prone to overfitting and inflexibility with new or noisy data. In contrast, probability-based models like Naive Bayes Xu (2018) and Hidden Markov Models Rabiner (1989) offered better generalization for tasks like spam detection and speech recognition, handling sequential and complex data effectively. Geometry-based approaches, including SVMs Joachims (2002) and dimensionality reduction techniques like PCA Wold et al. (1987), excelled in managing high-dimensional spaces but sometimes struggled with large datasets and computational efficiency. Finally, statistic-based methods like KNN Guo et al. (2003) and Logistic Regression Genkin et al. (2007) provided insights through nearest neighbor analysis and probability estimation, although they were sensitive to data preprocessing and could falter with nonlinear data.
II-B Deep Learning Approaches
Deep learning has become a key technology in text classification, capable of handling complex language features. CNN text classification models Kim (2014) capture local textual features through convolutional layers. LSTM Sutskever et al. (2014) and GRU Chung et al. (2014), as optimized versions of RNNs, are particularly effective in addressing long-distance dependencies in text. Zhou et al. (2024b) proposes an optical character recognition and classification method for cigarette laser code recognition, using a convolutional recurrent neural network to extract image features and utilizing BiLSTM for text classification. Transformer models, like BERT Devlin et al. (2018), achieve remarkable results in various NLP tasks by utilizing self-attention mechanisms. Specifically, the BERT model demonstrates powerful capabilities in text classification tasks. Raza et al. (2024) proposes the text recognition framework Nbias for detecting and eliminating biases, including data, corpus construction, model development, and evaluation layers. The dataset is collected from various fields, and a transformer-based token classification model is applied. Zhou et al. (2024a) proposes a semi-supervised generative adversarial learning method that improves the model’s classification performance with limited annotated data through generative adversarial networks. Jamshidi et al. (2024) introduces a hybrid model that combines BERT, LSTM, and Decision Templates (DT) for IMDB and Drug Review classification. However, these methods typically require substantial data for training, often necessitating extensive datasets to achieve optimal performance. This reliance on large training sets can pose challenges, especially when collecting or labeling data is difficult or impractical.
II-C LLM Approaches
LLMs represent a significant advancement in the field of text classification, building on the DL foundations that have revolutionized NLP. These models, which include notable examples such as GPT Brown et al. (2020), T5 Raffel et al. (2020), RWKV Peng et al. (2023), Mamba Gu & Dao (2023), Gemini Team et al. (2023), PaLM Chowdhery et al. (2023), Llama Touvron et al. (2023), and Claude Anthropic (2024), leverage massive amounts of data and extensive training regimes to understand and generate human-like text. Their ability to capture nuanced language patterns and context makes them highly effective for text classification tasks across various domains.
Recent studies have begun to explore the practical applications of LLMs in specialized fields. For instance, Chae & Davidson (2023) study the application of LLMs in sociological text classification, demonstrating their potential in social science research. Loukas et al. (2023) examine the performance and cost trade-offs when employing LLMs for text classification, focusing on financial intent detection datasets. Another study by Wei et al. (2023) investigates the effect of fine-tuning LLMs on text classification tasks within legal document review, highlighting how domain-specific adjustments can enhance model performance. Furthermore, the research identified as Cloutier & Japkowicz (2023) refines LLM performance on multi-class imbalanced text classification tasks through oversampling techniques, addressing one of the common challenges in ML.
Despite these advancements, there remains a gap between state-of-the-art methods and their accessibility to users without deep technical expertise or resources for extensive model training and tuning. The proposed smart expert system aims to bridge this gap by utilizing pre-trained LLMs as out-of-the-box classifiers that require minimal additional processing or domain-specific adaptation. This approach simplifies the deployment process and democratizes access to advanced NLP tools, enabling a wider range of users to benefit from cutting-edge technology without prohibitive costs or steep learning curves.
In summary, while traditional ML and NN methods have necessitated laborious preprocessing steps and dataset-specific model training, pre-trained LLMs offer an elegant solution that bypasses these requirements without compromising accuracy or effectiveness. By harnessing their inherent understanding of language nuances and context sensitivity, pre-trained LLMs can be directly applied to diverse text classification scenarios with remarkable efficiency—ushering in a new era where sophisticated NLP capabilities are readily available to all who seek them.
III Methodology
III-A Smart Expert System
Our proposed system integrates LLMs to refine the traditional text classification system, as illustrated in Figure 3. The framework of our expert system presents a comprehensive strategy that capitalizes on the strengths of LLMs while mitigating their traditional limitations.
Initially, our system aggregates data from many sources, either public or private. Unlike traditional ML/NN methods, which often require extensive retraining or fine-tuning when confronted with new data types, our LLM-based system can effectively adapt to these varied inputs without additional training. This versatility is one of the key strengths of our approach.
Subsequently, the system harnesses domain-specific data through zero-shot prompting or few-shot learning techniques or by fine-tuning a pre-trained LLM. This adaptive phase meticulously tailors the LLM’s capabilities to suit the target domain’s particular linguistic features and contextual subtleties, thereby bolstering accuracy and relevance for classification tasks.
Furthermore, the involvement of a domain knowledge expert is crucial but optional. They configure the system by establishing customized prompts that direct the LLM toward generating pertinent and contextually aware responses. This human-in-the-loop methodology guarantees that the system adheres to specific domain requirements and can adeptly manage intricate query scenarios.
Additionally, an LLM API serves as an intermediary between the model and user interface, enabling seamless real-time interactions. Through this user-friendly interface, users without expertise can effortlessly query the system for advice, classification results, sentiment assessments, predictions, or recommendations based on their input.
Lastly, our system incorporates an evaluation subsystem dedicated to continuously monitoring LLM performance. It scrutinizes accuracy and error rates while observing model behavior over time. Such vigilance facilitates perpetual enhancements and updates via model versioning and caching strategies.
By amalgamating these components, our Smart Expert System simplifies and elevates text classification processes in terms of adaptability and precision. It significantly diminishes reliance on domain experts for complex preprocessing or algorithmic design tasks—thereby democratizing access to cutting-edge NLP technologies across various sectors such as e-commerce and social media analytics.
III-B Evaluation metrics
We utilized several key metrics to assess the performance of LLMs as text classifiers. These metrics provide insights into the accuracy, precision, recall, and stability of the LLMs in handling classification tasks.
III-B1 Accuracy
This metric measures the proportion of correct predictions made by the model out of all predictions. It is calculated using the formula:
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
where $TP$ is the number of true positives, $TN$ is the number of true negatives, $FP$ is the number of false positives, $FN$ is the number of false negatives.
III-B2 F1 Score
The F1 score is a harmonic mean of precision and recall, providing a balance between them. It is particularly useful when dealing with imbalanced classes.
First, we define precision and recall:
$\text{Precision}=\frac{TP}{TP+FP}$ (2)
$\text{Recall}=\frac{TP}{TP+FN}$ (3)
Then, the F1 score is given by:
$F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
III-B3 U/E Rate
We propose a novel metric called Uncertainty/Error Rate (U/E rate) to evaluate the stability and reliability of LLM outputs. This metric quantifies the frequency at which an LLM either refuses to classify content or provides an output deemed unrelated or beyond its capabilities. The U/E rate is defined as:
[…]
where $U$ is the number of uncertain outputs (e.g., refusals to classify), $E$ is the number of erroneous outputs (e.g., unrelated or hallucinated results), and $N$ is the total number of test samples.
The U/E rate complements traditional performance metrics by highlighting instances where LLMs exhibit behavior divergent from deterministic ML/NN models, such as refusing to analyze content or producing hallucinated results.
By employing these evaluation metrics, we aim to provide a multifaceted view of LLM performance that encompasses traditional aspects like accuracy and F1 score and novel considerations introduced by their unique operational characteristics.
IV Dataset
The empirical evaluation of zero-shot text classification capabilities of LLMs in this study is conducted across four datasets, each chosen for its unique characteristics and relevance to specific domains. The datasets are as follows:
IV-A COVID-19-related Tweets Dataset
TABLE I: COVID-19-related Tweets Dataset Statistics
Negative Neutral Positive Total
Train 15398 7712 18046 41156
Test 1633 619 1546 3798
The first dataset consists of tweets related to the COVID-19 pandemic, curated by Preda (2020). This dataset is particularly valuable for sentiment analysis due to the real-time nature of Twitter data and the global impact of the pandemic, which has elicited a wide range of sentiments from users worldwide. As shown in Table I, it comprises a total of 41,156 training instances and 3,798 test instances, categorized into negative, neutral, and positive sentiments.
IV-B Economic Texts Dataset
TABLE II: Economic texts Dataset Statistics
Negative Neutral 2 Positive Total
Train 483 2302 1091 3876
Test 121 576 272 969
The second dataset includes economic texts compiled by Malo et al. (2014), designed for sentiment analysis within the financial domain. Economic texts present different linguistic features than social media content, often containing specialized vocabulary and complex expressions. The dataset contains 3,876 training samples and 969 test samples distributed across negative, neutral, and positive classes, as detailed in Table II. This dataset includes 5 levels of sentiment, which were merged into 3 levels in this study.
IV-C E-commerce Texts Dataset
TABLE III: E-commerce texts Dataset Statistics
Household Books C&A Electronics Total
Train 15449 9456 6936 8497 40338
Test 3863 2364 1734 2124 10085
For multi-class classification tasks beyond binary or ternary sentiment analysis, we utilize an e-commerce texts dataset provided by Gautam (2019). This dataset encompasses product reviews from four categories: Household items, Books, Clothing & Accessories (C&A), and Electronics. It offers a practical application scenario where accurate classification can directly influence recommendation systems and customer satisfaction. The training set includes 40,338 instances, while the test set contains 10,085 instances, as outlined in Table III.
IV-D SMS Spam Collection Dataset
TABLE IV: SMS Spam collection Statistics
Normal Spam Total
Train 3859 598 4457
Test 966 149 1115
Lastly, we incorporate an SMS Spam Collection dataset assembled by Almeida & Hidalgo (2012) to evaluate spam detection performance. Spam detection is critical for maintaining communication integrity and user trust in messaging platforms. This binary classification task involves distinguishing between normal messages and spam with a total of 4,457 training messages and 1,115 test messages presented in Table IV.
Each dataset was selected not only for its relevance to current societal issues but also because it collectively provides a comprehensive assessment environment that spans various text classification challenges faced by LLMs today. These include varying lengths of text inputs (from short tweets to longer reviews), domain-specific language usage (as seen in economic texts), diverse sentiment expressions (ranging from public health concerns to consumer products), and practical applications such as spam filtering. The diversity among these datasets will enable us to evaluate zero-shot learning capabilities across different contexts using LLMs robustly.
V Experimental Results
V-A Experiment Setup
Our experiment setup is designed to evaluate the performance of various models across different categories, ensuring a comprehensive analysis of the proposed methods. The models are categorized as follows:
  • Traditional ML Algorithms: This category includes MNB, LG, RF, DT, and KNN.
  • NN Architectures: We utilize advanced deep neural network models such as RNN, LSTM, and GRU.
  • ZSL Models: We explore zero-shot learning capabilities using transformer-based models, including BART (facebook/bart-large-mnli) and DeBERTa (microsoft/deberta-large-mnli).
  • LLMs: State-of-the-art LLMs including closed source models: GPT-3.5(gpt-3.5-turbo-0125), GPT-4 (gpt-4-1106-preview), Gemini-pro, and open source models: Llama3-8B(Llama3-8B-Instruct), Qwen-Chat(7B and 14B), and Vicuna-v1.5(7B and 13B) were assessed.
To maintain consistency in evaluation, the input processing was standardized for all traditional ML algorithms and NN architectures. Each model receives the same processed text derived from a uniform raw text processing pipeline applied to training and testing datasets. This standardization ensures that any observed variations in performance can be attributed more directly to the intrinsic capabilities of each model rather than disparities in input processing.
Conversely, unprocessed raw text from the testing dataset was used to fully leverage their natural language understanding abilities for zero-shot learning models and LLMs. It is important to note that this testing dataset remains consistent across all model types to provide a fair comparison platform.
In addition to these measures, we implemented a sampling strategy for dataset selection that respects the original label distribution within both training and test sets:
  • For datasets with more than 10,000 instances in their training set, only 10,000 instances were selected while preserving the original label distribution proportionally through stratified sampling.
  • Similarly, for test sets with more than 800 instances, only 800 instances were chosen based on their original label distribution.
This approach ensures that smaller datasets are fully represented while larger ones are sampled appropriately without introducing bias or altering their inherent class distributions.
Furthermore, when configuring prompts for LLMs within the experiments, uniformity is ensured by keeping prompts identical across different LLMs for the same dataset. When dealing with different datasets, a consistent core structure is maintained within prompts—only adjusting labels and dataset names as necessary—to minimize variability due to prompt differences.
By adopting these methodological choices—standardized input processing for traditional ML algorithms/NN architectures and careful instance selection respecting original distributions—we aim to provide an equitable experimental environment that allows us to assess each model’s inherent zero-shot text classification capabilities under comparable conditions.
V-B Experimental Results
The following tables present the experimental results for all the models. Notably, when employing few-shot strategies or fine-tuning, they are indicated by ”(S)” in the table or denoted by ”(F).” The value within the parenthesis shows the difference between the base LLM model result and the model employed with the S or F strategy.
TABLE V: COVID-19-RELATED TWEETS Sentiment classification results
Model ACC($\uparrow$) F1($\uparrow$) U/E($\downarrow$)
MNB 0.4037 0.3827 -
LR 0.3875 0.3131 -
RF 0.4462 0.3633 -
DT 0.4037 0.3416 -
KNN 0.3825 0.3481 -
GRU 0.6913 0.6324 -
LSTM 0.6687 0.6312 -
RNN 0.6600 0.6332 -
BART 0.5138 0.3638 -
DeBERTa 0.5375 0.3804 -
GPT-3.5 0.5550 0.5435 0.0000
GPT-4 0.5100 0.5054 0.0000
Gemini-pro 0.5025 0.5105 0.0388
Llama-3-8B 0.5112 0.5149 0.0013
Qwen-7B 0.4913 0.4689 0.0025
Qwen-14B 0.4562 0.4569 0.0100
Vicuna-7B 0.3600 0.3403 0.0000
Vicuna-13B 0.5050 0.4951 0.0013
Gemini-pro(S) 0.4888(-0.014) 0.4880(-0.022) 0.0375(-0.001)
Llama-3-8B(S) 0.5363(+0.025) 0.5298(+0.015) 0.0000(-0.001)
Qwen-7B(S) 0.3900(-0.101) 0.3519(-0.117) 0.0150(+0.012)
Qwen-14B(S) 0.4575(+0.001) 0.4556(-0.001) 0.0037(-0.006)
Vicuna-7B(S) 0.3700(+0.010) 0.3362(-0.004) 0.0013(+0.001)
Vicuna-13B(S) 0.5050(+0.000) 0.4951(+0.000) 0.0000(-0.001)
Llama-3-8B(F) 0.4675(-0.044) 0.4910(-0.024) 0.1175(+0.116)
Qwen-7B(F) 0.8388(+0.348) 0.8433(+0.374) 0.0000(+0.000)
S: with few shot strategy; F: with fine-tuned strategy
Table V presents results for the COVID-19-related text dataset. All models have relatively low performance except the fine-tuned LLM model of Qwen-7B, which performed the best in all metrics with $0.8388$ in accuracy and $0.8433$ F1 score and clearly provided all the answers.
Traditional algorithms, such as MNB, LR, RF, DT, and KNN, show poor accuracy and F1 scores, with RF performing marginally better than others in its class. In contrast, NN-based models demonstrate superior performance, with GRU leading in both ACC and F1 metrics. Among LLMs, before fine-tuning, GPT-3.5 exhibits the highest ACC and F1 scores, outperforming other LLMs, including GPT-4, while the performance is below NN methods. However, once the fine-tuning method was employed, the Qwen-7B(F) outperformed all the other models, including GRU’s best model. The best accuracy increased from 0.6913, performed by GRU, to 0.8388, and the F1 score from 0.63332, performed by RNN, to 0.8433.
TABLE VI: E-Commercial Product text classification results
Model ACC($\uparrow$) F1($\uparrow$) U/E($\downarrow$)
MNB 0.2562 0.2384 -
LR 0.3825 0.2873 -
RF 0.4875 0.3958 -
DT 0.4263 0.4165 -
KNN 0.3762 0.3414 -
GRU 0.9387 0.9383 -
LSTM 0.9363 0.9398 -
RNN 0.8975 0.9010 -
BART 0.7175 0.7246 -
DeBERTa 0.6025 0.6121 -
GPT-3.5 0.9125 0.9152 0.0063
GPT-4 0.9137 0.9221 0.0088
Gemini-pro 0.8775 0.8873 0.0100
Llama-3-8B 0.9113 0.9112 0.0000
Qwen-7B 0.5850 0.6584 0.1850
Qwen-14B 0.6575 0.6843 0.0800
Vicuna-7B 0.7100 0.7164 0.0050
Vicuna-13B 0.8363 0.8503 0.0138
Gemini-pro(S) 0.8862(+0.009) 0.8963(+0.009) 0.0100(+0.000)
Llama-3-8B(S) 0.9062(-0.005) 0.9065(-0.005) 0.0000(+0.000)
Qwen-7B(S) 0.6737(+0.089) 0.8226(+0.164) 0.1812(-0.004)
Qwen-14B(S) 0.7887(+0.131) 0.8548(+0.170) 0.0775(-0.003)
Vicuna-7B(S) 0.7925(+0.083) 0.7899(+0.074) 0.0000(-0.005)
Vicuna-13B(S) 0.9075(+0.071) 0.9153(+0.065) 0.0088(-0.005)
Llama-3-8B(F) 0.9175(+0.006) 0.9164(+0.003) 0.0000(+0.000)
Qwen-7B(F) 0.9713(+0.386) 0.9713(+0.313) 0.0000(-0.185)
S: with few shot strategy; F: with fine-tuned strategy
Table VI presents results for the e-commerce product text classification dataset. The GRU model shows the best performance among all models except for fine-tuned LLMs with an accuracy of $0.9387$ and an F1 score of $0.9383$, making it the leading model in these categories before considering fine-tuning. This illustrates the capability of GRU to handle sequence and context effectively, which is crucial for product text classification. Like Table V, traditional algorithms exhibit much lower accuracy and F1 scores than NN-based models. Among LLMs, GPT-based models also show impressive results before fine-tuning, with GPT-3.5 achieving slightly higher metrics than GPT-4. Applying fine-tuning techniques to LLMs such as Qwen-7B can result in superior accuracy of $0.9713$ and F1 scores of $0.9713$, making these models particularly effective for specialized tasks such as e-commerce product text classification.
TABLE VII: ECONOMIC TEXTS Sentiment Classification Results
Model ACC($\uparrow$) F1($\uparrow$) U/E($\downarrow$)
MNB 0.2600 0.2570 -
LR 0.5962 0.3055 -
RF 0.6375 0.4048 -
DT 0.4813 0.3805 -
KNN 0.5325 0.3528 -
GRU 0.6837 0.5494 0.7938
LSTM 0.6950 0.5967 0.7786
RNN 0.6550 0.4298 0.7754
BART 0.4125 0.4152 -
DeBERTa 0.4025 0.4119 -
GPT-3.5 0.6175 0.6063 0.0000
GPT-4 0.7638 0.7659 0.0000
Gemini-pro 0.7488 0.7519 0.0013
Llama-3-8B 0.7675 0.7710 0.0013
Qwen-7B 0.7550 0.7585 0.0025
Qwen-14B 0.7850 0.7860 0.0050
Vicuna-7B 0.7425 0.7250 0.0000
Vicuna-13B 0.6750 0.6735 0.0013
Gemini-pro(S) 0.6925(-0.056) 0.7217(-0.030) 0.0400(+0.039)
Llama-3-8B(S) 0.7550(-0.012) 0.7585(-0.013) 0.0013(+0.000)
Qwen-7B(S) 0.6837(-0.071) 0.6900(-0.069) 0.0288(+0.026)
Qwen-14B(S) 0.7738(-0.011) 0.7748(-0.011) 0.0063(+0.001)
Vicuna-7B(S) 0.7738(+0.031) 0.7607(+0.036) 0.0000(+0.000)
Vicuna-13B(S) 0.7575(+0.082) 0.7616(+0.088) 0.0013(+0.000)
Llama-3-8B 0.7913(+0.024) 0.7796(+0.009) 0.0000(-0.001)
Qwen-7B(F) 0.8400(+0.085) 0.8302(+0.074) 0.0000(-0.003)
S: with few shot strategy; F: with fine-tuned strategy
Table VII presents results for the economic texts sentiment classification dataset. The models show a broad performance spectrum, with the best results observed in fine-tuned LLMs. Traditional models continue to exhibit relatively low accuracy and F1 scores. RF performs somewhat better within this group but remains significantly lower than advanced models with a $0.6375$ accuracy and a $0.4048$ F1 score. NN-based models perform adequately, with GRU notably achieving an accuracy of $0.6837$ and an F1 score of $0.5494$. However, their performance is outstripped by more sophisticated models. LLM models, like GPT-4 and Gemini-pro, show significant improvements over traditional models, with GPT-4 reaching an accuracy of $0.7638$ and an F1 score of $0.7659$, indicating robust capabilities in processing complex economic texts. The fine-tuned models Llama-3-8B(F) and Qwen-7B(F) exhibit exceptional performance, with Qwen-7B(F) standing out for its remarkable accuracy and F1 score improvements. It is the only model that surpasses 80% accuracy and F1 score.
TABLE VIII: SMS SPAM COLLECTION Classification Results
Model ACC($\uparrow$) F1($\uparrow$) U/E($\downarrow$)
MNB 0.7488 0.6376 -
LR 0.8575 0.5419 -
RF 0.8962 0.7196 -
DT 0.8287 0.6559 -
KNN 0.8237 0.6241 -
GRU 0.9675 0.9257 -
LSTM 0.9675 0.9237 -
RNN 0.9725 0.9366 -
BART 0.7137 0.4943 -
DeBERTa 0.7025 0.5630 -
GPT-3.5 0.4988 0.5601 0.0000
GPT-4 0.9463 0.9495 0.0000
Gemini-pro 0.6500 0.7395 0.0575
Llama-3-8B 0.3937 0.4426 0.0025
Qwen-7B 0.7050 0.7527 0.0013
Qwen-14B 0.9137 0.9208 0.0000
Vicuna-7B 0.2762 0.2847 0.0000
Vicuna-13B 0.4550 0.5149 0.0000
Gemini-pro(S) 0.8163(+0.166) 0.8759(+0.136) 0.0488(-0.009)
Llama-3-8B(S) 0.5825(+0.189) 0.6482(+0.206) 0.0088(+0.006)
Qwen-7B(S) 0.7525(+0.047) 0.8124(+0.060) 0.0362(+0.035)
Qwen-14B(S) 0.8525(-0.061) 0.8730(-0.048) 0.0025(+0.003)
Vicuna-7B(S) 0.5675(+0.291) 0.6310(+0.346) 0.0013(+0.001)
Vicuna-13B(S) 0.6412(+0.186) 0.6976(+0.183) 0.0000(+0.000)
Llama-3-8B(F) 0.9825(+0.589) 0.9826(+0.540) 0.0000(-0.003)
Qwen-7B(F) 0.9938(+0.289) 0.9937(+0.241) 0.0000(+0.000)
S: with few shot strategy; F: with fine-tuned strategy
Table VIII details the SMS spam collection classification results, showcasing a notable disparity in model effectiveness, with fine-tuned LLMs and NN-based models outperforming others by a wide margin. Once again, traditional models underperform compared to NN and some LLM models, with RF leading the traditional pack but not nearly matching the performance of advanced models. NN-based models show exceptionally high performance, with RNN achieving the best results with an accuracy of $0.9725$ and an F1 score of $0.9366$. For LLMs, while some models demonstrate their high abilities in detecting spam SMS with accuracy high to more than 90%, like GPT-4 and Qwen-14B, some models failed in this task with accuracy lower than 0.5, like GPT-3.5, Llama-3-8B, and Vicuna families which are far worse than traditional ML methods or NN models. Notably, fine-tuning dramatically enhances the performance of models like Llama-3-8B(F) and Qwen-7B(F), which achieved the highest scores in both accuracy and F1 score with the values of $0.9938$ and $0.9927$, with the latter reaching near-perfect accuracy and F1 scores, highlighting the transformative power of model adaptation.
VI Discussion
VI-A Prompting strategy
The effectiveness of the few-shot strategy has been previously established; however, our investigation reveals that its influence is not uniform across different models and datasets.
In the context of Table V, five out of six models showed only marginal performance changes when employing this strategy. However, Qwen-7B(S) significantly underperformed with accuracy and F1 scores dropping by over 10%. This trend was not mirrored in Table VI, where four models marginally improved accuracy. Contrarily, Llama-3-8B(S) experienced a slight decrease, whereas Qwen-14B(S) notably excelled with an increase exceeding 13%. Table VII mostly saw marginal decreases in four out of six models, with only two showing minor improvements. These mixed results highlight that the impact of few-shot learning is highly model and dataset-dependent.
Table VIII, a different pattern emerged: while Qwen-7B(S) and Qwen-14B(S) underwent marginal changes in accuracy (4% increase and 6% decrease respectively), the other four models achieved significant improvements, Vicuna-7B(S), notably surged by over 25%. As for U/E metrics across datasets, there were minor variations except for specific trends within each dataset; COVID-19-related tweets fluctuated both ways, e-commercial product texts predominantly decreased or remained unchanged, while Spam SMS and economic texts mostly saw increases. These observations underscore that while few-shot strategies can be potent tools for model enhancement, their application requires careful consideration of the interplay between model architectures and dataset nuances to harness their potential fully.
VI-B Fine-tuning strategy
Our research involved fine-tuning two LLMs across four datasets as presented in Table V, VI, VII, and VIII, with the results indicating a significant enhancement in text classification performance. Notably, the Llama-3-8B(F) model did not show an improvement in the COVID-19-related tweets sentiment dataset, as presented in Table V. However, this model’s accuracy increased dramatically from 0.3937 to 0.9825 in spam SMS detection, transitioning from one of the least effective to one of the most proficient models, second only to Qwen-7B(F).
The Qwen-7B(F) model exhibited remarkable improvements across all datasets post-fine-tuning, with accuracy improved ranging from 0.085 to 0.386, thereby establishing it as a state-of-the-art model for these tasks. These findings highlight the potential of fine-tuning as a pivotal strategy for optimizing LLMs’ performance on specific text classification tasks.
More importantly, after fine-tuning, the U/E value cross models and datasets are down to 0, except for the Llama-3-8B(F) in tweet classification. This improvement in the standardized output makes the result consistent and makes the smart expert system more reliable.
Our results strongly advocate incorporating fine-tuning into LLM deployment workflows to unlock their full potential in specialized text classification scenarios.
VI-C Limitations
While LLMs demonstrated impressive proficiency in text classification across various domains and sentiment analysis, our experiments also uncovered a range of limitations that affect their practicality as zero-shot text classifiers. These constraints became particularly evident when juxtaposing LLMs with traditional ML algorithms and NN architectures.
  • Inconsistency in Standardized Output Formats: One key limitation of LLMs is their inconsistent output format. This can hinder the seamless integration of LLMs into existing systems that rely on specific output formats. For instance, when expecting results in a standardized JSON format, the LLMs may occasionally deviate by providing unstructured text strings or introducing variations in the keys or values within the JSON output. This inconsistency poses a challenge for downstream applications that require consistent and structured data for further processing.
  • Restrictions on Content Classification: Another limitation is the potential reluctance of certain LLMs to classify specific types of content. This reluctance can stem from various reasons, such as containing sensitive or impolite content. Furthermore, some LLMs may decline to classify content that exceeds their processing capabilities. Such limitations can constrain the scope of applications where LLMs can be effectively deployed, particularly in diverse or nuanced textual content scenarios.
  • Proprietary Model Constraints: Commercial closed-source LLMs introduce several operational constraints that can impede scalability and accessibility. API request rate limits service providers impose can throttle throughput and affect real-time application performance. Additionally, the cost associated with API calls can become prohibitive at scale, especially when dealing with large volumes of data requiring classification.
  • Hardware Resource Demands: The computational demands of utilizing LLMs, especially larger models, can impose significant hardware requirements. This includes the need for robust CPU and GPU resources to support the intensive processing capabilities of LLMs. The scalability and deployment of LLMs may be hindered by these hardware constraints, particularly in environments where access to high-performance computing resources is limited or costly. Detailed insights into the specific hardware prerequisites can offer valuable guidance for practitioners seeking to leverage LLMs effectively.
  • Time-intensive Processing: LLMs often exhibit longer inference times than traditional ML or NN models. The complex architecture and extensive parameterization of LLMs contribute to the prolonged processing duration required for generating classification results. This time-consuming nature can impact real-time or high-throughput applications requiring rapid responses, highlighting the trade-offs between accuracy and efficiency when employing LLMs for text classification tasks.
VII Conclusion and Future Work
In conclusion, our study has demonstrated the potential of LLMs as effective text classifiers within Smart Expert Systems, often surpassing traditional ML and NN approaches. Despite some LLMs showing lower accuracy in certain cases, strategic fine-tuning has proven to be a powerful method to enhance their domain-specific performance.
Our study underscores the transformative impact of LLMs on text classification, particularly in their ability to streamline the process by obviating the need for labor-intensive data preprocessing and bespoke model development. This characteristic is especially beneficial for small businesses seeking cost-effective solutions to integrate intelligent text classification into their operations without requiring deep expertise in ML or DL. The implications of our findings are profound: LLMs not only democratize access to cutting-edge AI technology but also empower organizations with limited resources to leverage sophisticated NLP tools. By harnessing the power of LLMs, businesses can efficiently process user feedback, enhance spam detection mechanisms, and automate workflows with minimal engineering effort.
For future work, we aim to address the challenges associated with fine-tuning, such as the need for substantial computational resources and expertise. Alternative strategies to improve LLM performance could include more detailed background provisions and precise label definitions. Refining non-standardized outputs remains a priority; employing a secondary LLM to process initial results could reduce user effort rates and improve overall system efficiency. As we continue refining these models, we aim to make LLM-powered expert systems more accessible and user-friendly across various applications.